40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10884     Entry time: Fri Jan 9 14:13:02 2015
Author: manasa 
Type: Update 
Category: General 
Subject: ALS noise at ~30Hz 

Before starting to move things around the PSL table, I measured the ALS out of loop noise for reference. The noise spectrum showed excess noise around 30Hz.  The traces in blue were taken before I went in and those in red were taken after my work.

I also looked at the power spectrum of LSC-TRY_OUT to see if the noise is from the IR lock itself and there was nothing suspicious with it. 

Since the green transmission was 50% lower than the usual, I touched the steering mirrors for green at the Y end table to get a better transmission and beat signal. The noise still hasn't disappeared. 

I am not sure if this could be from our neigbors who have been running rock experiments since morning. We should check back to see if this noise disappears once they shut down.

DTT xml files are available in directory /users/manasa/data/150109/

Attachment 1: ALSYoutLoop.png  71 kB  | Hide | Hide all
ALSYoutLoop.png
Attachment 2: LSC_TRY.png  63 kB  | Hide | Hide all
LSC_TRY.png
ELOG V3.1.3-